# Clarifying Ambiguities: on the Role of Ambiguity Types in Prompting Methods for Clarification Generation

### v1

@@ 6. Task 2: Information Retrieval (IR)
This section aims to investigate the impact of integrating ambiguity into LLM reasoning on IR performance. Table 8 shows IR results of the two different interaction modes (select + respond) and the baseline without clarification. We detail the result alongside three successive turns. Generally, we observe that AT-CoT $>$ CoT $>$ AT-standard $\approx$ standard for most of the interaction modes and turns. For instance, clarifications obtained with the method AT-COT allow to reach the best IR metrics values for the TREC Web Track 2013-2014 dataset over all turns (0.397, 0.391, and 0.381 for each turn respectively vs 0.392, 0.384, and 0.373 at most for the baselines). We also note that IR performance is always better for the respond interaction mode corresponding to the generation of clarifying questions (in contrast to the select mode based on query reformulation. Altogether, these results highlight two main conclusions: 1) it aligns with our remarks on the CG performance, demonstrating the benefits of introducing ambiguity-oriented LLM reasoning for clarification, both intrinsically and extrinsically. And 2) this reinforces our hypothesis based on the need for clarification interactions based on ambiguity and reasoning in IR. Our findings also demonstrate the robustness of our methodology in interaction scenarios. For both interaction scenarios select and respond, AT-CoT consistently provides the best IR performance, implying that our method can adapt to various real-world scenarios such as query suggestion-based scenarios (e.g. search suggestions) or chat scenarios (e.g. chatbot).

### v2

@@ 6. Task 2: Information Retrieval (IR)
This section aims to investigate the impact of integrating ambiguity into LLM reasoning on IR performance. Table 8 shows IR results of the two different interaction scenarios (select & respond) and the baseline without clarification. We detail the result alongside three successive turns. Generally, we observe that AT-CoT $>$ CoT $>$ AT-standard $\approx$ standard for most of the interaction modes and turns. For instance, clarifications obtained with the method AT-COT allow to reach the best IR metrics values for the TREC Web Track 2013-2014 dataset over all turns (0.397, 0.391, and 0.381 for each turn respectively vs 0.392, 0.384, and 0.373 at most for the baselines). We also note that IR performance is always better for the respond interaction mode corresponding to the generation of clarifying questions (in contrast to the select mode based on query reformulation. Altogether, these results highlight two main conclusions: 1) it aligns with our remarks on the CG performance, demonstrating the benefits of introducing ambiguity-oriented LLM reasoning for clarification, both intrinsically and extrinsically. And 2) this reinforces our hypothesis based on the need for clarification interactions based on ambiguity and reasoning in IR. Our findings also demonstrate the robustness of our methodology in interaction scenarios. For both interaction scenarios select and respond, AT-CoT consistently provides the best IR performance, implying that our method can adapt to various real-world scenarios such as query suggestion-based scenarios (e.g. search suggestions) or chat scenarios (e.g. chatbot).